# A Sensitivity Analysis Framework for Quantifying Confidence in Decisions in the Presence of Data Uncertainty

### v1

@@ 3.4 Step 4: Assess confidence in decision
Having constructed the CID metric, we can investigate the analyst’s confidence in their reference decision (to intervene) when data are MNAR. We do so over ranges of $t$ for the two missingness mechanisms introduced previously. Figure 3(a) and Figure 3(b) display the results. These figures also display graphically for five values of $t$ the multiple imputation point estimates of the empirical frequencies of the lead levels among the $N$ children. The decision changes around $t=0.4$ under the accordion-style missingness mechanism and around $t=0.8$ under the parametric-style missingness mechanism. As shown in Table 2, for the accordion-style mechanism, a value of $t=0.5$ generates an estimate of 19% of children having lead levels above three, which is below the decision-rule threshold. For the parametric-style missingness mechanism, a value of $t=1$ results in a similar 19% figure, though the distribution is somewhat more concentrated at levels of 2 and 3. For larger values of $t$, the costs of overestimation increase with departures from MAR.
@@
To evaluate the plausibility of such departures from MAR, we turn to reference lead levels among the population of children in the United States published by the Centers for Disease Control (CDC), as reported by Feldman et al. (2024). In particular, the CDC estimates that around half of U.S. children have levels less than or equal to 1, around 75% have levels less than or equal to 2, and around 90% have levels less than or equal to 3. To approach these U.S.  percentages, we need very large values of $t$. Although lead levels among children in North Carolina may not match those in the CDC estimates exactly, it seems not unrealistic for $t$ to take on values large enough to change the decision about intervention, resulting in a decision under MAR that could incur relatively large cost under (11). Because of this, we recommend that the analyst should not have confidence in the decision to intervene under the MAR analysis.
4 Discussion
@@
Our examples used analyses with limited numbers of variables. In these cases, departures from the assumptions used in the reference analysis can be encoded with single parameters. It may be possible to define univariate knobs in contexts with richer analyses. For example, sensitivity analyses in causal inference of observational studies often posit a single unmeasured confounder. In some contexts, however, the analyst may seek to turn multiple knobs corresponding to multiple assumptions. Developing interpretable and practical methods for multivariate confidence in decision analysis is an important topic for future research.

### v2

@@ 3.4 Step 4: Assess confidence in decision
Having constructed the CID metric, we can investigate the analyst’s confidence in their reference decision (to intervene) when data are MNAR. We do so over ranges of $t$ for the two missingness mechanisms introduced previously. Figure 3(a) and Figure 3(b) display the results. These figures also display for five values of $t$ the multiple imputation point estimates of the empirical frequencies of the lead levels among the $N$ children. The decision changes around $t=0.4$ under the accordion-style missingness mechanism and around $t=0.8$ under the parametric-style missingness mechanism. As shown in Table 2, for the accordion-style mechanism, a value of $t=0.5$ generates an estimate of 19% of children having lead levels above three, which is below the decision-rule threshold. For the parametric-style missingness mechanism, a value of $t=1$ results in a similar 19% figure, though the distribution is somewhat more concentrated at levels of 2 and 3. The cost of this overestimation is visualized as $05/.54\approx.10$, which is incurred for all values of $t$ that cause $D_{t}=0$. The figures also reveal large costs that accrue quickly for $t<0$, which would occur if, for some reason, in actuality more than 25% of children have lead levels exceeding 3.
@@
To evaluate the plausibility of such departures from MAR, we turn to reference lead levels among the population of children in the United States published by the Centers for Disease Control (CDC), as reported by Feldman et al. (2025+). In particular, the CDC estimates that around half of U.S. children have levels less than or equal to 1, around 75% have levels less than or equal to 2, and around 90% have levels less than or equal to 3. If the population lead levels in North Carolina in fact are similar to these U.S. percentages, we would have values of $t$ that represent substantial selection bias. For example, with the parametric missingness mechanism, a value of $t\geq 10$ results in completed-data estimates
@@
near the CDC percentages. However, as evident in Table 2, the CID analysis indicates that the decision to intervene changes at much smaller values of $t$ than $t=10$. Further, the reference decision about intervention could incur relatively large costs in the CID metric. For example, if we set $t=2$, we have a true percentage of children with lead levels exceeding 3 of 15%. Thus, by using the MAR estimate of 25%, if $t=2$ we would be deciding to intervene unnecessarily, incurring around 10% of the maximum possible cost of an incorrect decision based on $\hat{\theta}=0.25$. Translating this to a more practicable scale, if the state would spend, say, \$10 million per percentage point above the true population percentage, they would be spending an additional \$50 million when the reality might not dictate such intervention. Because of the decision change at values of $t$ that seem plausible, and because we would incur a substantial cost at those values, we recommend that the analyst should not have confidence in the decision to intervene under the MAR analysis.
4 Further Considerations for Specifying CID Metric
As noted previously, the specification of the CID metric should be tailored to the analysis and decision at hand. Nonetheless, analysts can follow certain principles to assist development of the CID metrics, as we now describe.
We consider two parts to constructing a CID metric. The first part concerns encoding the decision rule, as done, for example, in (2) and (5). To establish actionable and realistic decision rules, analysts can work with domain experts or policy-makers. In general, we expect them to be binary or ternary functions.
The second part concerns encoding the cost function that is incurred from inaccuracies in the estimate of interest. This specification is akin to specifying a cost (or utility) function, as done in decision analysis and in economics applications (Chung, 1994; Coto-Millán, 2003). In our applications, we use linear functions to represent costs of inaccuracies in analyses. In the Bread and Peace analysis, for example, this linear function implies that (i) analysts consider larger confidence interval overlaps to be more desirable than smaller confidence interval overlaps and (ii) analysts consider the marginal cost of the same decrease in the overlap to be constant. Alternatively, analysts could specify a nonlinear function that expresses the costs of inaccuracies, for example, one that encodes a relatively flat cost when overlap exceeds 0.5 and drops sharply thereafter. We expect that, in many cases, analysts will want to enforce monotonicity when setting the cost function.
If analysts assign a constant cost function, effectively they just use the decision rule as the CID metric. For example, in the Bread and Peace analysis, the CID metric could be based purely on whether or not the prediction interval includes 0.5, irregardless of the confidence interval overlap. In this case, the CID framework encourages analysts to consider whether the ultimate decision—not just the estimates—changes when assessing sensitivity to data imperfections. As noted previously, a 5% change in a point estimate from a reference analysis can be interpreted differently depending on how close the point estimate is to the boundary of a decision change.
Some decision rules also can be viewed through the lens of significance testing. For example, in the North Carolina lead data analysis, with a constant cost function one could perform a significance test of the null hypothesis $\theta<0.20$ based on the multiple imputations for a given posited missingness distribution. However, in other contexts significance testing may not be as natural to implement, for example, with a three-option decision task like the one in the Bread and Peace CID analysis.
In economics applications and decision theory, analysts often use loss (utility) functions to select from competing strategies. The analyst selects the strategy that optimizes the loss function, according to some posited distribution of potential outcomes. Our sensitivity analysis context differs. First, if the analyst puts highest probability on a certain assumptions about the data imperfections, for example, a certain measurement error or missingness mechanism, it is rational for them to use that assumptions for the reference analysis. The resulting CID metric would provide information on confidence in decisions for that reference analysis; optimization is not involved. That said, given a probability distribution on the data imperfections, e.g., on the values of $t$ in our examples, it should be possible to compute an expected CID value. This expectation can serve as a one-number summary of the confidence in decision metric for the reference analysis.
5 Discussion
@@
Our examples used analyses with limited numbers of variables. In these cases, departures from the assumptions used in the reference analysis can be encoded with single parameters. It may be possible to define univariate knobs in contexts with richer analyses. For example, sensitivity analyses in causal inference of observational studies often posit a single unmeasured confounder. Starting from an analysis with no unmeasured confounding (the reference analysis), they determine how strong the confounder needs to be to make the estimated causal effect vanish (turning a knob until the decision changes). This is a CID analysis, which potentially could be enhanced with a metric that offers more nuance on the decisions, e.g., if different effect sizes result in different actions. In some contexts, however, the analyst may seek to turn multiple knobs corresponding to multiple assumptions. Developing interpretable and practical methods for multivariate confidence in decision analysis is an important topic for future research.